f(x)=4+3^x
Question: 1. State the inverse of the function
2. State the domain of f(x) and f^-1(x)
I'm SoOOOOOOOOO LOST HELP!!!!!!!!!!

Answer by stanbon(75887)

f(x)=4+3^x
Question: 1. State the inverse of the function
Interchange x and y to get:
x=4+3^y
Solve for y:
3^y=x-4
y=log(base3)(x-4)
--------------
2. State the domain of f(x) and f^-1(x)
Domain of f(x) is all Real Numbers
Domain of f^-1(x) is all Real Numbers >4
=============
Cheers,
Stan H.